OpenStudy (anonymous):

find the value of the expression if sin beta= 0.45, find cos(pi/2-beta)

OpenStudy (anonymous):

do i just do pi/2-0.45?

OpenStudy (lgbasallote):

sin beta is 0.45...but beta is not 0.45 does that make sense?

OpenStudy (lgbasallote):

\[\Large \cos (\frac \pi 2 - \beta) \implies \cos \frac \pi 2 \cos \beta + \sin \frac \pi 2 \sin \beta\] are you familiar with this?

OpenStudy (anonymous):

no idk what you just did

OpenStudy (lgbasallote):

it's a trig formula \[\LARGE \cos (A - B ) \implies \cos A \cos B + \sin A \sin B\]

OpenStudy (lgbasallote):

in this case, our A is pi/2 and B is beta make sense?

OpenStudy (anonymous):

yeah but whats beta?

OpenStudy (anonymous):